One month, Rehna used 1/6 of her monthly salary for a car payment. She also paid rent, which was 25% more than the car payment. What fraction of her monthly salary did Rehana use that month for the car payment and rent combined?
A. 5/24 B. 3/8 C. 5/12 D. 12

Found 3 solutions by stanbon, jim_thompson5910, milo1954:
Answer by stanbon(75887)

One month, Rehna used 1/6 of her monthly salary for a car payment. She also paid rent, which was 25% more than the car payment. What fraction of her monthly salary did Rehana use that month for the car payment and rent combined?
-------
Ans: (1/6) + (1/4) = (2/12) + (3/12) = 5/12
==================
A. 5/24 B. 3/8 C. 5/12 D. 12
========
Cheers,
Stan H.
=========

Answer by jim_thompson5910(35256)

25% of 1/6 = 0.25 * (1/6) = (1/4)*(1/6) = 1/24

Add this to 1/6 to get 1/6 + 1/24 = 4/24 + 1/24 = 5/24

So the fraction going to rent is 5/24

The fraction going to the car payment is 1/6 (given)


Add the two fractions: 1/6 + 5/24 = 4/24+5/24 = 9/24 = 3/8

=======================================================

Final Answer: 3/8

Answer by milo1954(10)

car payment=C= 1/6 S=S/6
rent =R= 1.25 C
C+R=(1/6)S+(1.25/6)S
C+R=(2.25/6)S
C+R==(225/600)S
C+R=(3/8)S